OpenStudy (lilai3):

@robtobey and @AccessDenied please help! 2.2 yd cubed equals to how many ft cubed?

OpenStudy (anonymous):

1 cu yard = 3^3 = 27 cubic feet\[2.2 \text{ yd}^3 = 2.2 * 27 \text{ ft}^3=59.4 \text{ ft}^3 \]

OpenStudy (lilai3):

so your answer is 59.4 ft square? or is it a cubiced

OpenStudy (lilai3):

i'm sorry but can you explain the steps you did and why you did it?

OpenStudy (anonymous):

It is 59.4 cubic ft

OpenStudy (lilai3):

can you explain the steps you did and why you did it?

OpenStudy (anonymous):

A cubic yard is equal to 27 cubic feet. If you have 2.2 of them then you have 2.2 * 27 cubic feet or 59.4 cubic feet.

OpenStudy (lilai3):

oh i get it now thank you

OpenStudy (anonymous):

You're welcome and thank you for the medal.